OpenStudy (amber15):

A private plane traveled from Seattle to the rugged wilderness at a average speed of 176 mph. on the return trip the average speed was 308 mph. if the total traveling time was 5 hours. How far is Seattle to the wilderness? please help

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

guess you need to use distance equals rate times time right?

OpenStudy (misty1212):

time equal distance divided by rate

OpenStudy (amber15):

yes

OpenStudy (amber15):

i'm lost on this bad

OpenStudy (misty1212):

time going is \[\frac{D}{175}\] and time returning is \[\frac{D}{308}\] the total time is \[\frac{D}{175}+\frac{D}{308}\] which you know is \(5\)

OpenStudy (amber15):

yes

OpenStudy (misty1212):

oops i made a mistake, should be 176 no matter

OpenStudy (amber15):

right

OpenStudy (misty1212):

solve \[\frac{D}{176}+\frac{D}{308}=5\] for \(D\)

OpenStudy (amber15):

ok

OpenStudy (misty1212):

it is not fun but i am sure you can do it

OpenStudy (amber15):

560

OpenStudy (misty1212):

let me know when you get it

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (amber15):

so its 560 correct

OpenStudy (misty1212):

yes

OpenStudy (amber15):

thank you

OpenStudy (misty1212):

you are welcome

OpenStudy (amber15):

can you help me with one more